### v1

# Structured dataset of reported cloud seeding activities in the United States (2000–2025) using a large language model

Abstract
We present a structured dataset of reported cloud seeding activities in the U.S. from 2000 to 2025, including the year, season, state, seeding agent, apparatus used for deployment, and purpose. Using OpenAI’s o4-mini large language model (LLM), combined with multi-stage PDF-to-text conversion and response-parsing code, we processed 836 historical reports from the National Oceanic and Atmospheric Administration (NOAA) to extract the data. The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo. Our results address the gap in accessible cloud seeding data in the United States, and demonstrate the potential of LLMs to extract structured environmental data from historical documents. More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
Background & Summary
Cloud seeding is a weather modification technique used to enhance precipitation, typically in regions affected by drought, water scarcity, or to support snowpack accumulation.
In the United States, cloud seeding has been practiced since the 1940s but reporting was not standardized until the Weather Modification Reporting Act of 1972 [1]. The Act requires individuals and organizations to notify the U.S. Department of Commerce at least 10 days before and after conducting weather modification operations. These reports are submitted to NOAA via email and compiled on the NOAA Weather Modification website [2].
Currently, NOAA’s publicly available records on their website are available for 2000-2025. We expected earlier records to be available and submitted a data request to NOAA for documents covering 1972–1999, but have not received a response to date. As a result, the dataset currently covers 2000-2025. If older cloud seeding reports from 1972–1999 are recovered and made publicly available, the same pipeline can be applied with minimal modifications to expand the dataset’s coverage.
While the NOAA reports are made publicly available online, they are stored individually as scanned PDFs with inconsistent formatting and structure. This lack of standardization makes it difficult to extract key information at scale, or integrate the reports into other structured datasets. As a result, data on cloud seeding in the U.S., such as locations, dates, purposes, cloud seeding agent, and agent deployment methods, remains largely inaccessible for analysis.
To solve this data access problem, we created a structured dataset from 836 NOAA cloud seeding reports spanning from 2020 to 2025. Our work transforms these scattered, inconsistently formatted documents into a single, structured CSV file that enables the analysis of U.S. weather modification activities over time. The dataset has 94.72% accuracy on all extracted fields when compared with the original PDFs.
This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management—many of which exist in PDF form with highly variable formatting. By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest.
Beyond its methodological contributions, the resulting dataset has potential value across a wide range of applications. Researchers can use it to study long-term patterns in weather modification practices, analyze the use and evolution of different seeding agents and deployment methods, or assess geographic and seasonal trends in operational activity.
For example, while there have been numerous recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g. the 2017 SNOWIE campaign in Western Idaho[3]), there has been little qualitative analysis of reported historical operational cloud seeding activities. With recent renewed interest in both traditional weather modification technology (e.g. the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geo-engineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change, this dataset can provide historical context for weather modification usage in the US. In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of weather modification usage[4].
Methods
Figure 1: NOAA Data Extraction Pipeline using LLM
Input Data Source
Cloud seeding reports were manually downloaded from the NOAA Central Library’s Weather Modification Project Reports archive [2]. A total of 1,025 PDF files were downloaded, covering all reported weather modification activities in the United States from 2000-2025.
Each PDF contains the initial report on the planned cloud seeding project (Form 17-4), a second tabular report describing interim and final details (Form 17-4A), or a combination of both. Form 17-4 contains project-level summaries, including information on the purpose of the project (e.g., precipitation enhancement or hail suppression), the operational period, and the agents and apparatus used. Form 17-4A provides more granular operational details such as the number of seeding hours, the type and amount of agent released, and meteorological conditions on specific dates.
Beyond these two forms, many reports include additional supporting materials such as operational maps, meteorological analyses, auxilliary tables, cloud seeding decision logs, and narrative summaries. While these materials provide valuable context for understanding cloud seeding activities, they also increase variability across reports, requiring additional preprocessing and file handling.
Preprocessing
A custom Python pipeline was developed to organize and preprocess the PDF documents (Figure 1). First, all downloaded files were stored in a local directory. In many cases, a single cloud seeding project was documented across multiple submissions, indicated by suffixes such as .M (initial report), .I (interim report), and .F (final report). Since these reports refer to the same geographic area and operational period, they were programmatically consolidated to preserve project-level information and ensure completeness during data extraction. After consolidation, there were 836 unique weather modification projects to process and include in our dataset.
The scanned documents included both typed and hand-written text (and sometimes a mix of both in the same document). To convert the merged PDFs into plain text, three approaches were used. For native, text-based documents (57% of documents), text was extracted using the pymupdf Python library [5]. For scanned or hand-written documents (43% of documents), optical character recognition (OCR) was applied using a combination of pytesseract, an open-source wrapper for the Tesseract OCR engine [6] (processed 16% of all documents), and llm-whisperer (processed 27% of all documents), a layout-preserving PDF extraction technology which uses both OCR and text-based extraction [7]. All code for the preprocessing pipeline, including file merging and text extraction, is available in a public GitHub repository [8].
For each document, the text from Form 17-4 was extracted using a three-tie approach: first attempting pymupdf, then pytesseract, and lastly llm-whisperer.
Only Form 17-4 was used because it contained all the key metadata we needed for our dataset, including: purpose, date ranges, description of agent and apparatus used, and the target area. Truncating the input in this way also helped reduce noise and minimize costs. In addition to the extracted text, the original filename was also provided to the model because the filenames often contained useful contextual clues about the report’s location or time frame (e.g. North Dakota Cloud Modification Project - District I_02-1147_06.01.2002-08.31.2002.pdf).
Notably, there appeared to be two distinct naming conventions in the archive. Older files (generally before 2017) tend to use a descriptive format that includes the location and date (e.g., Colorado-SanJuan-2003.pdf), whereas newer files (from around 2018 onward) follow a standardized alphanumeric format such as 2018TXSWMA-1_2018.pdf, which embeds the year, state abbreviation, and project code. These filenames were leveraged during preprocessing and extraction to assist with project disambiguation and metadata recovery.
LLM Integration
After experimenting with various model options, we decided to use OpenAI’s o4-mini model due to its balance of performance and cost [9]. We also experimented with three different prompt templates, and the final prompt was chosen to use chain-of-thought reasoning with example reasoning steps to extract fields aligned with Form 17-4 and return structured text outputs for subsequent parsing. All prompt versions and evaluation scripts are available in the GitHub repository [8].
LLM Prompt
# NOAA Weather Modification Report Extraction Expert
You are an expert data extractor specialized in parsing historical NOAA Weather Modification Activity reports. For each report, utilize the PDF-converted text and filename to extract 9 critical fields.
## Instructions
For each field, carefully analyze all available information using a step-by-step reasoning process. Clearly document your reasoning, resolve conflicting or ambiguous information logically, and then provide your final extracted value using the Final Extracted Fields Format.
## Explanation of Fields
The fields you must extract are detailed below, with guidelines for each:
... truncated for brevity ...
## Example Chain-of-Thought Reasoning
**YEAR OF WEATHER MODIFICATION ACTIVITY:**
- Filename: ‘2018UTNORT-1.pdf
- Filename starts with ‘2018‘, clearly indicating the year.
- No conflicting year mentioned elsewhere.
- Final inference: **2018**
... truncated for brevity ...
[Continue similarly structured reasoning for all fields]
## Final Extracted Fields Format
... truncated for brevity ...
Postprocessing
The LLM response was a list containing the key information we requested to extract:
YEAR: [extracted value]
SEASON: [extracted value]
STATE: [extracted value]
AGENT: [extracted value]
APPARATUS: [extracted value]
PURPOSE: [extracted value]
TARGET AREA: [extracted value]
START DATE: [extracted value]
END DATE: [extracted value]
Then, in order to neatly store this information into CSV format, we mapped key values to more concise column names, stripped any blank space, and converted to lower case. The resulting data object now looked like this:
parsed_gpt_response = {’start_date’: ’10/15/2007’, end_date’: ’12/24/2007’, season’: winter’, target_area’: alta and snowbird ski resorts’, year’: ’2007’, state’: utah’, type_of_agent’: silver iodide’, type_of_apparatus’: ground’, purpose’: augment snowpack’, filename’: Alta - Snowbird, Utah_07-1384_10.15.2007-12.24.2007.pdf’}
With the formatted data object, we could easily save the data to CSV with each object as a separate row in the CSV, and each key-value pair corresponding to the columns.
Data Records
The final cloud seeding dataset is provided as a single CSV file titled cloud_seeding_us_2000_2025.csv, using one row per cloud seeding project and one column per extracted field. It is publicly available on Zenodo at the following DOI: https://doi.org/10.5281/zenodo.14925812. It contains structured information extracted from 836 cloud seeding project reports in the United States between 2000 and 2025.
Figure 2: Dataset Screenshot
The following fields are present in the dataset:
  • filename: The original NOAA report filename from which the data was extracted (e.g., 2021TXSWMA_STWMA_2021.pdf).
  • year: The calendar year in which the cloud seeding activity occurred (e.g., 2015).
  • season: The meteorological or operational season during which the activity took place (e.g., winter, summer).
  • state: The U.S. state where the cloud seeding operation was conducted (e.g., california)
  • agent: The material used to seed clouds (e.g., silver iodide, dry ice).
  • apparatus: The method or equipment used to deliver the seeding agent (e.g., ground, airborne).
  • purpose: The stated objective of the cloud seeding activity (e.g., augment snowpack, increase rain).
  • target_area: The geographic area targeted for cloud seeding (e.g., lake tahoe and truckee river basin).
  • start_date: The reported start date of the cloud seeding activity (e.g., 11/15/2014).
  • end_date: The reported end date of the cloud seeding activity (e.g., 4/15/2015)
Data Visualizations
Between 2000 and 2025, most U.S. cloud seeding activity occurred in Utah, Colorado, and eastern California, primarily aimed at enhancing snowpack through ground-based silver iodide dispersion. This pattern is consistent with the nine western states known to have operational cloud seeding programs: California, Colorado, Idaho, Nevada, New Mexico, North Dakota, Texas, Utah, and Wyoming [4]. In Texas, by contrast, efforts focus on increasing rainfall and suppressing hail using a combination of airborne silver iodide dispersion and ground-based shockwave generation. Overall, cloud seeding activity declined between 2002 and 2023, but showed an uptick in 2024, with 2025 also on track for increased activity.
Figure 3: U.S. Cloud Seeding Activity by State (2000-2025)
Figure 4: Cloud Seeding Activity over Time (2000-2025)
Figure 5: Purpose of Cloud Seeding Activity
Technical Validation
A key challenge in using LLMs for data extraction is their generative nature. Although LLMs are capable of producing fluent and structured text, they can generate information that is not present in the source material. In this context, such hallucinated content may include incorrect years, locations, or project descriptions, which could compromise the reliability of downstream analyses involving historical weather modification activities in the United States.
Additional difficulties arise from the quality of the input documents. Many PDFs contain handwritten annotations, low-resolution scans, or incomplete text. For example, some entries refer the reader to information from a previous year or external attachments rather than providing complete answers in the designated fields. These issues reduce the clarity and completeness of the source data, which in turn affects extraction accuracy.
A common source of error involved missing form data. In several cases, including reports from the North Dakota Cloud Modification Project and the Western Kansas Weather Modification Program, submitters left fields blank and instead referred to separate documents such as an operational plan or an attached sheet. These references could not be processed directly, limiting the ability of the model to extract a full set of structured values.
To assess and improve the accuracy of the extracted dataset, we conducted a structured technical validation. This involved comparing the model’s outputs against a set of manually annotated ground truth records, allowing us to quantify field-level accuracy and identify common failure cases across the dataset.
Accuracy Evaluations
To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM. The review was manually conducted independently by two human experts. A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level. For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-ground-truth.py.
Table 1: Field-Level Accuracy (n=200)
Field Accuracy (%)
Year 100.00%
Season 87.94%
State 100.00%
Agent 89.95%
Apparatus 92.96%
Purpose 97.49%
Target Area not annotated
Start Date not annotated
End Date not annotated
Overall Average 94.72%
Model Comparisons
We experimented with various OpenAI models to improve accuracy, including: gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3. This allowed us to observe how accuracy changed with model complexity and specialization. Ultimately, OpenAI’s ”more affordable reasoning model”, o4-mini, achieved the best balance of performance and cost. When comparing the different model variants, the same prompt was used (Prompt C).
Table 2: Model Comparison (n=50) using Prompt C
LLM Accuracy (%)
gpt-4.1 88.33%
gpt-4.1-mini 83.67%
gpt-4o-mini 86.67%
o3 90.33%
o4-mini 88.00%
Prompt Engineering
We also evaluated how different prompts influenced extraction accuracy, holding the model constant with gpt-4.1-mini. We compared three prompt formats: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and chain-of-thought (Prompt C) that encouraged step-by-step reasoning. Prompt C, using chain-of-thought reasoning, substantially outperformed the others, achieving an accuracy of 87.67%. This suggests that explicit reasoning instructions with examples can improve structured data extraction accuracy.
Table 3: Prompt Comparison (n=50) using gpt-4.1-mini
Prompt Accuracy (%)
Prompt A (Detailed) 72.00%
Prompt B (Concise) 57.00%
Prompt C (Chain-of-Thought) 87.67%
Together, these validation efforts demonstrate that LLM-based data extraction can be improved with careful model selection and prompt design. However, while we tested several variants, we did not exhaustively test every permutation of models-and-prompts listed. Further experimentation might reveal additional insights.
Usage Notes
All weather modification reports are self-reported by project operators and may not represent a complete record of weather modification activity (e.g., if there are omissions). For PDF-to-text conversion, note that llm-whisperer is free only for up to 100 PDF page extractions per day. Additionally, note that the OpenAI models referenced (e.g., gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3), incur usage costs depending on model and volume, so users should consult current OpenAI pricing for budgeting larger-scale replication [10].
Code Availability
Python code for extracting the dataset from the NOAA reports can be obtained from the public repository at https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor. Python code for analyzing and visualizing the dataset is available at: https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor.
Acknowledgments
Both authors acknowledge funding from Columbia University’s Data Science Institute. K.D.L. acknowledges funding from the Zegar Family Foundation.