### v2

# Structured dataset of reported cloud seeding activities in the United States (2000–2025) using a large language model

Abstract
Cloud seeding, a weather modification technique aimed at locally enhancing precipitation, has been used in the Western United States since the 1940’s. However, datasets to facilitate large-scale analysis of cloud seeding efforts are not currently available. To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S. from 2000 to 2025, including the year, season, state, seeding agent, apparatus used for deployment, and purpose. Using OpenAI’s o4-mini large language model (LLM), combined with multi-stage PDF-to-text conversion and response-parsing code, we processed 836 historical reports from the National Oceanic and Atmospheric Administration (NOAA) to extract the data. The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo. Our results address the gap in accessible cloud seeding data in the United States, and demonstrate the potential for LLMs to extract structured environmental data from historical documents. More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
Background & Summary
Cloud seeding is a weather modification technique used to enhance precipitation, typically in regions affected by drought, water scarcity, or to support snowpack accumulation.
In the United States, cloud seeding has been practiced since the 1940s but reporting was not standardized until the Weather Modification Reporting Act of 1972 [1]. The Act requires individuals and organizations to notify the U.S. Department of Commerce at least 10 days before and after conducting weather modification operations. These reports are submitted to NOAA via email and compiled on the NOAA Weather Modification website [2].
While the NOAA reports are made publicly available online, they are stored individually as scanned PDFs with inconsistent formatting and structure. This lack of standardization makes it difficult to extract key information at scale, or integrate the reports into other structured datasets. As a result, data on cloud seeding in the U.S., such as locations, dates, purposes, cloud seeding agent, and agent deployment methods, remains largely inaccessible for analysis.
To solve this data access problem, we created a structured dataset from 836 NOAA cloud seeding reports spanning from 2000 to 2025. Our work transforms these scattered, inconsistently formatted documents into a single, structured CSV file that enables the analysis of U.S. weather modification activities over time. The dataset has 94.72% accuracy on all extracted fields when compared with the original PDFs.
This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management—many of which exist in PDF form with highly variable formatting. By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest[3].
Beyond its methodological contributions, the resulting dataset has potential value across a wide range of applications. Researchers can use it to study long-term patterns in weather modification practices, analyze the use and evolution of different seeding agents and deployment methods, or assess geographic and seasonal trends in operational activity.
For example, while there have been several recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g. the 2017 Seeded and Natural Orographic Wintertime clouds: the Idaho Experiment (SNOWIE) campaign in Western Idaho[4]), there has been little qualitative analysis of reported historical operational cloud seeding activities. With recent renewed interest in both traditional weather modification technology (e.g. the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geo-engineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change[5], this dataset can provide historical context for weather modification usage in the US. In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of large-scale weather modification[3].
Methods
Figure 1: NOAA Data Extraction Pipeline
Input Data Source
Cloud seeding reports were manually downloaded from the NOAA Central Library’s Weather Modification Project Reports archive [2]. A total of 1,025 PDF files were downloaded, covering all reported weather modification activities in the United States from 2000-2025.
Each PDF contains the initial report on the planned cloud seeding project (Form 17-4), a second tabular report describing interim and final details (Form 17-4A), or a combination of both. Form 17-4 contains project-level summaries, including information on the purpose of the project (e.g., precipitation enhancement or hail suppression), the operational period, and the agents and apparatus used. Form 17-4A provides more granular operational details such as the number of seeding hours, the type and amount of agent released, and meteorological conditions on specific dates.
Beyond these two forms, many reports include additional supporting materials such as operational maps, meteorological analyses, auxilliary tables, cloud seeding decision logs, and narrative summaries. While these materials provide valuable context for understanding cloud seeding activities, they also increase variability across reports, requiring additional preprocessing and file handling.
Preprocessing
A custom Python pipeline was developed to organize and preprocess the PDF documents (Figure 1). First, all downloaded files were stored in a local directory. In many cases, a single cloud seeding project was documented across multiple submissions, indicated by suffixes such as .M (initial report), .I (interim report), and .F (final report). Since these reports refer to the same geographic area and operational period, they were programmatically consolidated to preserve project-level information and ensure completeness during data extraction. After consolidation, there were 836 unique weather modification projects to process and include in our dataset.
The scanned documents included both typed and hand-written text (and sometimes a mix of both in the same document). To convert the merged PDFs into plain text, three approaches were used. For native, text-based documents (57% of documents), text was extracted using the pymupdf Python library [6]. For scanned or hand-written documents (43% of documents), optical character recognition (OCR) was applied using a combination of pytesseract, an open-source wrapper for the Tesseract OCR engine [7] (processed 16% of all documents), and llm-whisperer (processed 27% of all documents), a layout-preserving PDF extraction technology which uses both OCR and text-based extraction [8]. All code for the preprocessing pipeline, including file merging and text extraction, is available in a public GitHub repository [9].
For each document, the text from Form 17-4 was extracted using a three-tie approach: first attempting pymupdf, then pytesseract, and lastly llm-whisperer.
Only Form 17-4 was used because it contained all the key metadata we needed for our dataset, including: purpose, date ranges, description of agent and apparatus used, and the target area. Truncating the input in this way also helped reduce noise and minimize costs. In addition to the extracted text, the original filename was also provided to the model because the filenames often contained useful contextual clues about the report’s location or time frame (e.g. North Dakota Cloud Modification Project - District I_02-1147_06.01.2002-08.31.2002.pdf).
Notably, there appeared to be two distinct naming conventions in the archive. Older files (generally before 2017) tend to use a descriptive format that includes the location and date (e.g., Colorado-SanJuan-2003.pdf), whereas newer files (from around 2018 onward) follow a standardized alphanumeric format such as 2018TXSWMA-1_2018.pdf, which embeds the year, state abbreviation, and project code. These filenames were leveraged during preprocessing and extraction to assist with project disambiguation and metadata recovery.
LLM Integration
After experimenting with various model options, we decided to use OpenAI’s o4-mini model due to its balance of performance and cost [10]. We also experimented with three different prompt templates, and the final prompt was chosen to use chain-of-thought reasoning with example reasoning steps to extract fields aligned with Form 17-4 and return structured text outputs for subsequent parsing (shown below). Chain-of-thought reasoning guides the model to break down complex tasks into intermediate steps, which improves performance on tasks requiring multi-step inference [11].
All prompt versions and evaluation scripts are available in the GitHub repository [9].
LLM Prompt
# NOAA Weather Modification Report Extraction Expert
You are an expert data extractor specialized in parsing historical NOAA Weather Modification Activity reports. For each report, utilize the PDF-converted text and filename to extract 9 critical fields.
## Instructions
For each field, carefully analyze all available information using a step-by-step reasoning process. Clearly document your reasoning, resolve conflicting or ambiguous information logically, and then provide your final extracted value using the Final Extracted Fields Format.
## Explanation of Fields
The fields you must extract are detailed below, with guidelines for each:
... truncated for brevity ...
## Example Chain-of-Thought Reasoning
**YEAR OF WEATHER MODIFICATION ACTIVITY:**
- Filename: ‘2018UTNORT-1.pdf
- Filename starts with ‘2018‘, clearly indicating the year.
- No conflicting year mentioned elsewhere.
- Final inference: **2018**
... truncated for brevity ...
[Continue similarly structured reasoning for all fields]
## Final Extracted Fields Format
Present your final extracted results concisely as follows, in lowercase, comma-separating multiple values when applicable.
Do not include commentary, explanations, or placeholders. Leave fields blank if truly unknowable after exhausting all inference methods.
... truncated for brevity ...
Postprocessing
The LLM response was a list containing the key information we requested to extract:
YEAR: [extracted value]
SEASON: [extracted value]
STATE: [extracted value]
AGENT: [extracted value]
APPARATUS: [extracted value]
PURPOSE: [extracted value]
TARGET AREA: [extracted value]
START DATE: [extracted value]
END DATE: [extracted value]
Then, in order to neatly store this information into CSV format, we mapped key values to more concise column names, stripped any blank space, and converted to lower case. The resulting data object now looked like this:
parsed_gpt_response = {’start_date’: ’10/15/2007’, end_date’: ’12/24/2007’, season’: winter’, target_area’: alta and snowbird ski resorts’, year’: ’2007’, state’: utah’, type_of_agent’: silver iodide’, type_of_apparatus’: ground’, purpose’: augment snowpack’, filename’: Alta - Snowbird, Utah_07-1384_10.15.2007-12.24.2007.pdf’}
With the formatted data object, we could easily save the data to CSV with each object as a separate row in the CSV, and each key-value pair corresponding to the columns.
Data Records
The final cloud seeding dataset is provided as a single CSV file titled cloud_seeding_us_2000_2025.csv, using one row per cloud seeding project and one column per extracted field. It is publicly available on Zenodo at the following DOI: https://doi.org/10.5281/zenodo.14925812. It contains structured information extracted from 836 cloud seeding project reports in the United States between 2000 and 2025.
Figure 2: Dataset Screenshot
The following fields are present in the dataset:
  • filename: The original NOAA report filename from which the data was extracted (e.g., 2021TXSWMA_STWMA_2021.pdf).
  • year: The calendar year in which the cloud seeding activity occurred (e.g., 2015).
  • season: The meteorological or operational season during which the activity took place (e.g., winter, summer).
  • state: The U.S. state where the cloud seeding operation was conducted (e.g., California)
  • agent: The material used to seed clouds (e.g., silver iodide, dry ice).
  • apparatus: The method or equipment used to deliver the seeding agent (e.g., ground, airborne).
  • purpose: The stated objective of the cloud seeding activity (e.g., augment snowpack, increase rain).
  • target_area: The geographic area targeted for cloud seeding (e.g., Lake Tahoe and Truckee River Basin).
  • start_date: The reported start date of the cloud seeding activity (e.g., 11/15/2014).
  • end_date: The reported end date of the cloud seeding activity (e.g., 4/15/2015)
Data Visualizations
Between 2000 and 2025, cloud seeding activity in the United States was geographically concentrated in a few key states (Figure 3). California, Colorado, Utah, and Texas accounted for the majority of operations, with activity highest in western states that rely on snowpack for water supply. These records align with states that maintain active cloud seeding programs.
Over time, the number of weather modification events peaked in the early to mid-2000s, then declined through the 2010s, before rebounding in 2024 and 2025 (Figure 4). This timeline shows that while cloud seeding has remained a consistent tool, its usage has fluctuated, likely due to policy changes, drought conditions, and funding cycles.
The primary stated purpose of cloud seeding events was to increase snowpack, followed by increasing precipitation and rain (Figure 5). Snowpack enhancement is particularly important in mountainous regions where runoff from melting snow supplies reservoirs and irrigation systems.
Finally, the types of agents and apparatus used vary significantly (Figure 6). Silver iodide is by far the most common seeding agent, especially in ground-based and airborne operations. Other agents such as calcium chloride and carbon dioxide are used less frequently. Ground deployment is the most prevalent method, particularly when using silver iodide, although a mix of airborne and dual (ground+airborne) approaches is also present.
Figure 3: U.S. Cloud Seeding Activity by State (2000–2025). States with active weather modification programs show the highest number of recorded operations.
Figure 4: Cloud Seeding Activity by State over Time (2000–2025). The number of activities peaked around 2005, declined gradually, and rose again after 2023.
Figure 5: Purpose of Cloud Seeding Activity (2000–2025). Snowpack enhancement is the leading purpose, followed by increasing precipitation and rain.
Figure 6: Agent and Apparatus Used for Cloud Seeding (2000–2025). Silver iodide dominates among agents and is deployed most frequently using ground-based apparatus.
Technical Validation
A key challenge in using LLMs for data extraction is their generative nature. Although LLMs are capable of producing fluent and structured text, they can generate information that is not present in the source material [12]. In this context, such hallucinated content may include incorrect years, locations, or project descriptions, which could compromise the reliability of downstream analyses involving historical weather modification activities in the United States.
Additional difficulties arise from the quality of the input documents. Many PDFs contain handwritten annotations, low-resolution scans, or incomplete text. For example, some entries refer the reader to information from a previous year or external attachments rather than providing complete answers in the designated fields. These issues reduce the clarity and completeness of the source data, which in turn affects extraction accuracy.
A common source of error involved missing form data. In several cases, including reports from the North Dakota Cloud Modification Project and the Western Kansas Weather Modification Program, submitters left fields blank and instead referred to separate documents such as an operational plan or an attached sheet. These references could not be processed directly, limiting the ability of the model to extract a full set of structured values.
To assess and improve the accuracy of the extracted dataset, we conducted a structured technical validation. This involved comparing the model’s outputs against a set of manually annotated ground truth records, allowing us to quantify field-level accuracy and identify common failure cases across the dataset.
Accuracy Evaluations
To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM. The review was manually conducted independently by two human experts. A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level. For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-golden.py.
Table 1: Field-Level Accuracy (n=200) using o4-mini LLM
Field Accuracy (%)
Year 100.00%
Season 87.94%
State 100.00%
Agent 89.95%
Apparatus 92.96%
Purpose 97.49%
Target Area not annotated
Start Date not annotated
End Date not annotated
Overall Average Accuracy 94.72%
Model Comparisons
We evaluated several OpenAI models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3 [10]. This comparison allowed us to assess how accuracy varied with model complexity and specialization. Among these, o4-mini offered the best balance of performance and cost, achieving 95.00% accuracy at approximately \$0.005 per document. While o3 achieved the highest accuracy at 96.33%, it was an order of magnitude more expensive, costing around \$0.05 per document.
Each accuracy value in Table 2 represents the average field-level accuracy across 50 annotated NOAA reports, using a consistent prompt (Prompt C). For each report, model outputs were compared to human-labeled values across all fields, and the proportion of correct fields was averaged across the dataset. Although large language models are known to generate outputs probabilistically, we observed stable performance across repeated runs of the same model.
Table 2: Model Comparison (n=50)
LLM Overall Average Accuracy (%)
gpt-4.1 93.33%
gpt-4.1-mini 93.33%
gpt-4o-mini 91.67%
o3 96.33%
o4-mini 95.00%
Prompt Engineering
We also evaluated how different prompts influenced extraction accuracy, holding the model constant (o4-mini). Three prompt formats were tested: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and a chain-of-thought format (Prompt C) that encouraged step-by-step reasoning. Prompt C achieved the highest accuracy at 95.00%, suggesting that explicit reasoning steps can substantially improve structured data extraction.
Prompt B performed poorly (44.00%) due to frequent missing values in the outputs, likely caused by its lack of field-level guidance and examples. In contrast, Prompt A performed better (93.67%) and benefitted from including worked examples for each field. While not explicitly labeled as chain-of-thought, Prompt A was influenced by the same principles, demonstrating that even partial integration of reasoning structure can enhance accuracy.
Table 3: Prompt Comparison (n=50) using o4-mini LLM
Prompt Overall Average Accuracy (%)
Prompt A (Detailed) 93.67%
Prompt B (Concise) 44.00%
Prompt C (Chain-of-Thought) 95.00%
Together, these validation efforts demonstrate that LLM-based data extraction can be significantly improved with thoughtful prompt engineering. Chain-of-thought reasoning and example-driven prompts, in particular, are well-suited for scientific data extraction from complex archival documents. While our evaluation covered key variants, further experimentation could uncover additional insights.
Usage Notes
All weather modification reports in this dataset are self-reported by project operators and may not represent a complete record of U.S. weather modification activity, particularly in cases of unsubmitted or omitted filings.
As of this writing, NOAA’s publicly available records span the years 2000–2025. We anticipated earlier records to be accessible and have submitted a formal data request to NOAA for archived documents covering 1972–1999. If these older reports are recovered and released, the current extraction pipeline can be applied with minimal modification to extend the dataset’s temporal coverage.
For PDF-to-text conversion, note that llm-whisperer is free only for up to 100 PDF page extractions per day. Additionally, note that the OpenAI models referenced (e.g., gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3), incur usage costs depending on model and volume, so users should consult current OpenAI pricing for budgeting larger-scale replication [13].
Code Availability
Python code for extracting the dataset from the NOAA reports can be obtained from the public repository at https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor. Python code for analyzing and visualizing the dataset is available at: https://github.com/jdonohue44/US-Cloud-Seeding-Analysis.
Acknowledgments
Both authors acknowledge funding from Columbia University’s Data Science Institute. K.D.L. acknowledges funding from the Zegar Family Foundation.